OpenStudy (anonymous):

During a clothing store’s Bargain Days, the regular price for T-shirts is discounted by $5. There is a state sales tax of 5%, and the $5 discount is applied before the sales tax is calculated. a. Write an expression that shows the regular price r of a T-shirt minus the $5 discount. b. Write a rule for the function p(r) that expresses the final price p of a T-shirt with the discount applied and sales tax added. c. How much would you pay during Bargain Days for a shirt regularly priced at $15.50?

OpenStudy (anonymous):

I did it like this. A.) t=r-5 B.) p(r)= r-5*0.525 C.) $11.03 Would this be correct??

OpenStudy (nykoreyan):

correct

OpenStudy (anonymous):

Yay! Thanks. I asked someone else and they really confused me. They said a) d = r - 5 where t is the ful price b) p = ( r - 5 ) * 1.05 c) p = (15.5-5)*1.05 = 10.5 * 1.05 = $ 11.02

OpenStudy (anonymous):

I don't even know where they got 1.05 from

OpenStudy (anonymous):

correct buddy

OpenStudy (anonymous):

ok thanks again.

OpenStudy (anonymous):

no problem

OpenStudy (nykoreyan):

y didn't I get the best response I told u it was right first lmao

OpenStudy (anonymous):

Sorry my computer kicked me out. That wasn't me though who gave the best response to him. I gave it to you. Thanks.